OpenStudy (anonymous):

What is the 43rd term of an arithmetic sequence with a rate of increase of -6 and a11 = 12? A. -174 B. -176 C. -180 D.-186 E. -240

hartnn (hartnn):

so, d=-6 a11= 12 firstly,use the formula \(\large a_n=a_1+(n-1)d\) to get the value of a1 (n = 11)

OpenStudy (anonymous):

A little confused still

hartnn (hartnn):

\(a_n=a_1+(n-1)d\) is the general formula we are given, a11 =12 ....so,n=11 d=-6 just plug these in and find the value of a1

OpenStudy (amistre64):

just let a11 = a1 a(11-10) = a1 a(43-10) = a33

OpenStudy (anonymous):

Isn't the answer -174?

OpenStudy (amistre64):

or in this case it might be even simpler to go to a0 an = 12 - 6n; a(43-11) = a32 = 12 -6(32)

OpenStudy (anonymous):

The answeris A?

OpenStudy (amistre64):

of course not ...

OpenStudy (amistre64):

12 - 6(32)

hartnn (hartnn):

how did u get -174 ? its actually 42nd term...

OpenStudy (anonymous):

180! Yes? And i'm in summer school. And my math teacher isnt here, so when my science teacher helped me I guess he was wrong..

OpenStudy (amistre64):

logically: 43 - 11 = 32 so you are 32 interation away. 32(-6) must be added to 12

hartnn (hartnn):

-180 yes, but more important is that you know exactly how to get it...

OpenStudy (amistre64):

*iterations ... my fingers hate me

OpenStudy (anonymous):

Thank you both.

hartnn (hartnn):

welcome ^_^

OpenStudy (amistre64):

good luck ;)